verify that each given function is a solution of the

differential equation.

5.y′′\:−\:y\:=\:0;\:y\left(t\right)\:=\:e^t


1
Expert's answer
2021-09-01T23:54:03-0400

Let us verify that the function "y(t)=e^t" is a solution of the differential equation "y''-y=0". Since "y'(t)=e^t" and "y''(t)=e^t," we conclude that "y''(t)-y(t)=e^t-e^t=0," and hence "y(t)=e^t" is a solution of the differential equation "y''-y=0".